# Hard-Constrained Neural Networks with Universal Approximation Guarantees

### v1

@@ 4.1 HardNet-Aff: Imposing Input-Dependent Affine Constraints
where $A(x)\in\mathbb{R}^{n_{\text{ineq}}\times n_{\text{out}}}$, $b(x)\in\mathbb{R}^{n_{\text{ineq}}}$, $C(x)\in\mathbb{R}^{n_{\text{eq}}\times n_{\text{out}}}$, $d(x)\in\mathbb{R}^{n_{\text{eq}}}$ for $n_{\text{ineq}}$ ineqaulity and $n_{\text{eq}}$ equality constraints. For partitions $A(x)=[A_{(:n_{\text{eq}})}\;A_{(n_{\text{eq}}:)}]$ and $C(x)=[C_{(:n_{\text{eq}})}\;C_{(n_{\text{eq}}:)}]$, we make the following assumptions about the constraints:
@@
i) The constraints in (4) are feasible,  ii) $C_{(:n_{\text{eq}})}$ is invertible for all $x\in\mathcal{X}$,
iii) $\tilde{A}(x):=A_{(n_{\text{eq}}:)}-A_{(:n_{\text{eq}})}C_{(:n_{\text{eq}})}^{-1}C_{(n_{\text{eq}}:)}$ has full row rank for all $x\in\mathcal{X}$.
Given $x\in\mathcal{X}$, when $C(x)$ has full row rank (i.e., no redundant constraints), there exists an invertible submatrix of $C(x)$ with its $n_{\text{eq}}$ columns. Without loss of generality, we can assume $C_{(:n_{\text{eq}})}$ is such submatrix by considering a proper permutation of the components of $f$. Then, the second assumption holds when the same permutation lets $C_{(:n_{\text{eq}})}$ invertible for all $x\in\mathcal{X}$. The last assumption requires the total number of the constraints $n_{\text{ineq}}+n_{\text{eq}}$ to be less than or equal to the output dimension $n_{\text{out}}$ and $A(x)$ to have full row rank.
Under the assumptions, we first efficiently reduce the $n_{\text{ineq}}+n_{\text{eq}}$ constraints to $n_{\text{ineq}}$ equivalent inequality constraints on partial outputs $f_{(n_{\text{eq}}:)}$ for a partition of the function $f(x)^{\top}=[f_{(:n_{\text{eq}})}^{\top}\;f_{(n_{\text{eq}}:)}^{\top}]$. Consider the hyperplane in the codomain $\mathcal{Y}$ over which the equality constraints are satisfied. Then, for the function output $f(x)$ to be on the hyperplane, the first part $f_{(:n_{\text{eq}})}$ is determined by $f_{(n_{\text{eq}}:)}$:
@@
and $A^{+}:=A^{\top}(AA^{\top})^{-1}$ is the pseudoinverse of $A$. This novel projection satisfies the following properties; see Appendix A.1 for a proof.
@@
where $\bar{f_{\theta}}(x):=\big{[}[C_{(:n_{\text{eq}})}^{-1}\big{(}d(x)-C_{(n_{\text{eq}}:)}f_{\theta}(x)\big{)}]^{\top}\;f_{\theta}(x)^{\top}\big{]}^{\top}\in\mathbb{R}^{n_{\textup{out}}}$.
Note that HardNet-Aff in (7) for $(n_{\text{ineq}},n_{\text{eq}})=(1,0)$ is equivalent to the single-inequality-constraint case in (3). However, unlike (3), which is the closed-form solution of the optimization in (2), HardNet-Aff, in general, does not perform the minimum $\ell^{2}$-norm projection for the constraint in (4). Instead, we can understand its projection geometrically based on Proposition 4.2 (iii). First, $\bar{f}_{\theta}(x)$ determined by $f_{\theta}(x)$ is on the hyperplane over which the equality constraints hold. Then, the projection preserves the distance from the boundary of the feasible set for each constraint when $\bar{f}_{\theta}(x)$ satisfies the constraint. Otherwise, the projected output is located on the boundary. Thus, HardNet-Aff projects the augmented output $\bar{f}_{\theta}(x)$ onto the feasible set in the direction parallel to the boundaries of the satisfied constraints’ feasible sets as described in Fig. 1.
While HardNet-Aff is guaranteed to satisfy the hard constraints in (4), it should not lose the neural network’s expressivity to be useful for deployment in practical applications. To that end, we provide a rigorous argument for HardNet-Aff preserving the neural network’s expressive power by the following universal approximation theorem; see Appendix A.2 for a proof.
@@ 4.2 HardNet-Cvx: Imposing General Input-Dependent Convex Constraints
Going beyond the affine constraints, we present HardNet-Cvx as a framework that enforces general input-dependent convex constraints:
$f(x)\in\mathcal{C}(x)\;\;\forall x\in\mathcal{X}$ (8)
where $\mathcal{C}(x)\subset\mathbb{R}^{n_{\text{out}}}$ is a convex set. Unlike the affine constraints case, we cannot extend the closed-form projection of the single constraint case in (3) for general convex constraints. Thus, we present HardNet-Cvx by generalizing the optimization-based projection in (2) as below:
@@
While no general closed-form solution for this optimization problem exists, we can employ differentiable convex optimization solvers such as Amos & Kolter (2017) for affine constraints (when their number exceeds the output dimension $n_{\text{out}}$) and Agrawal et al. (2019) for more general convex constraints. Originally, these solvers were introduced as new layer architectures within neural networks to enable learning optimization processes in latent spaces, rather than solving given optimization problems directly. In our approach, we freeze these layers and exploit their differentiability to backpropagate through the solutions. This allows us to project the output $f_{\theta}(x)$ onto the feasible set $\mathcal{C}(x)$ and train the projected function via conventional gradient-based algorithms. It is worth noting that this idea was mentioned in passing by Donti et al. (2021b) but was not pursued due to computational complexity concerns.
@@
Consider input-dependent constrained sets $\mathcal{C}(x)\subset\mathbb{R}^{n_{\textup{out}}}$ that are convex for all $x\in\mathcal{X}\subset\mathbb{R}^{n_{\textup{in}}}$. For any $p\in[1,\infty)$, let =$\mathcal{F}=\{f\in L^{p}(\mathcal{X},\mathbb{R}^{n_{\textup{out}}})|f(x)\in\mathcal{C}(x)\;\forall x\in\mathcal{X}\}$. Then, HardNet-Cvx with $w$-width ReLU neural networks defined in (9) universally approximates $\mathcal{F}$ if $w\geq\max\{n_{\textup{in}}+1,n_{\textup{out}}\}$.
@@ 5 Experiments
In this section, we demonstrate the versatility and effectiveness of HardNet over four scenarios with required hard constraints: fitting functions under constraints, learning optimization solvers, optimizing control policies in safety-critical systems, and learning safe decision logic for aircraft systems.
As evaluation metrics, we measure the violation of constraints in addition to the application-specific performance metrics. For a test sample $x\in\mathcal{X}$ and $n_{\text{ineq}}$ inequality constraints $C_{\leq}\big{(}f(x)\big{)}\leq 0\in\mathbb{R}^{n_{\text{ineq}}}$, their violation is measured with the maximum ($\leq$ max) and mean ($\leq$ mean) of $\mathsf{ReLU}\big{(}C_{\leq}\big{(}f(x)\big{)}\big{)}$ and the number of violated constraints ($\leq$ #). Similar quantities of $|C_{=}\big{(}f(x)\big{)}|$ are measured for $n_{\text{eq}}$ equality constraints $C_{=}\big{(}f(x)\big{)}=0\in\mathbb{R}^{n_{\text{eq}}}$. Then, they are averaged over all test samples. The inference time for the test set is also compared.
We compare HardNet with the following baselines: (i) NN: Plain neural networks, (ii) Soft: Soft-constrained neural networks. To penalize constraint violation, for a sample point $x_{s}\in\mathcal{X}$, additional regularization terms $\lambda_{\leq}\|\mathsf{ReLU}\big{(}C_{\leq}\big{(}f(x_{s})\big{)}\big{)}\|_{2}^{2}+\lambda_{=}\|C_{=}\big{(}f(x_{s})\big{)}\|_{2}^{2}$ are added to the loss function, (iii) NN+Proj/Soft+Proj: The projection of HardNet is applied at test time to the outputs of NN/Soft, (iv) DC3 (Donti et al., 2021b): Similarly to HardNet-Aff, DC3 takes a neural network that approximates the part of the target function. It first augments the neural network output to satisfy the equality constraints. Then, it corrects the augmented output to minimize the violation of the inequality constraints via the gradient descent algorithm. DC3 backpropagates through this iterative correction procedure to train the model. For all methods, we use 3-layer fully connected neural networks with 200 neurons in each hidden layer and $\mathsf{ReLU}$ activation function.
Figure 2: Learned functions at the initial (left) and final (right) epochs for the function fitting experiment. The models are trained on the samples indicated with circles, with their RMSE distances from the true function shown in parentheses. HardNet-Aff adheres to the constraints from the start of the training and generalizes better than the baselines on unseen data. On the other hand, the baselines violate the constraints throughout the training.
Table 2: Results for the function fitting experiment. HardNet-Aff generalizes better than the baselines with the smallest RMSE distance from the true function without any constraint violation. The max, mean, and the number of constraint violations are computed out of 401 test samples. Violations are highlighted in red. Standard deviations over 5 runs are shown in parentheses.
@@
NN 1.06 (0.21) 2.39 (0.42) 0.44 (0.12) 206.00 (9.27) 0.16 (0.01)
Soft 1.45 (0.26) 3.36 (0.62) 0.66 (0.16) 206.00 (9.44) 0.15 (0.00)
DC3 1.01 (0.11) 2.20 (0.39) 0.39 (0.06) 200.80 (6.34) 8.92 (1.51)
HardNet-Aff 0.40 (0.01) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00) 1.06 (0.24)
5.1 Function Fitting under Constraints
In this experiment, we demonstrate the efficacy of HardNet-Aff on a problem involving fitting a continuous function $f:[-2,2]\rightarrow\mathbb{R}$ shown in Fig. 2. The function outputs are required to avoid specific regions defined over separate subsets of the domain $[-2,2]$. These constraints can be expressed as a single input-dependent affine constraint on the function output. The models are trained on 50 labeled data points randomly sampled from $[-1.2,1.2]$; see Appendix A.4 for details.
As shown in Fig. 2 and Table 2, HardNet-Aff consistently satisfies the hard constraints throughout training and achieves better generalization than the baselines, which violate these constraints. Especially at the boundaries $x\!=\!-1$ and $x=1$ in the initial epoch results, the jumps in DC3’s output value, caused by DC3’s correction process, insufficiently reduce the constraint violations. The performance of its iterative correction process heavily depends on the number of gradient descent steps and the step size, DC3 requires careful hyperparameter tuning unlike HardNet-Aff.
@@ 5.2 Learning Optimization Solver
Table 3: Results for learning solvers of nonconvex optimization problems with 100 variables, 50 equality constraints, and 50 inequality constraints. HardNet-Aff attains feasible solutions with the smallest suboptimality gap among the feasible methods. The max, mean, and the number of violations are computed out of the 50 constraints. Violations are highlighted in red. Standard deviations over 5 runs are shown in parentheses.
@@
Optimizer -14.28 (0.00) 0.00/0.00/0 (0.00/0.00/0) 0.00/0.00/0 (0.00/0.00/0) 1019.3 (10.3)
NN -27.36 (0.14) 13.19/1.33/13.00 (2.28/0.43/1.99) 17.48/6.79/50 (4.72/0.80/0) 0.21 (0.01)
NN+Proj 768.46 (0.70) 0.00/0.00/0 (0.00/0.00/0) 0.00/0.00/0 (0.00/0.00/0) 8.99 (4.77)
Soft -12.62 (0.09) 0.07/0.00/0.31 (0.01/0.00/0.02) 0.31/0.10/49.97 (0.03/0.01/0.01) 0.23 (0.04)
Soft+Proj -12.18 (0.24) 0.00/0.00/0 (0.00/0.00/0) 0.00/0.00/0 (0.00/0.00/0) 18.57 (3.51)
DC3 -14.06 (0.03) 0.27/0.01/1.88 (0.04/0.00/0.29) 0.00/0.00/0 (0.00/0.00/0) 4.57 (0.08)
HardNet-Aff -13.78 (0.02) 0.00/0.00/0 (0.00/0.00/0) 0.00/0.00/0 (0.00/0.00/0) 6.74 (0.90)
We consider the problem of learning optimization solvers as in Donti et al. (2021b) with the following nonconvex optimization problem:
$f(x)=\operatorname*{arg\,min}_{y}\;\;\frac{1}{2}y^{\top}Qy+p^{\top}\sin{y}\;\;\text{ s.t. }\;Ay\leq b,\;Cy=x,$ (10)
@@
In this experiment, we guarantee that the given constraints are feasible for all $x\in[-1,1]^{n_{\text{eq}}}$ by computing a proper $b$ for randomly generated $A,C$ as described in Donti et al. (2021b). Then the models are trained on 10000 unlabeled data points uniformly sampled from $[-1,1]^{n_{\text{eq}}}$. To perform this unsupervised learning task, the loss function for each sample $x_{s}$ is set as $\frac{1}{2}f_{\theta}(x_{s})^{\top}Qf_{\theta}(x_{s})+p^{\top}\sin{f_{\theta}(x_{s})}$. As shown in Table 3, HardNet-Aff consistently finds feasible solutions with small suboptimality gap from the optimizer (IPOPT) with much shorter inference time.
5.3 Optimizing Safe Control Policy
Table 4: Results for optimizing safe control policies. HardNet-Aff generates trajectories that have zero constraint violation and have the smallest costs among the methods that have zero violation. The max and mean constraint violations are computed for the violations accumulated throughout the trajectories. Violations are highlighted in red. Standard deviations over 5 runs are shown in parentheses.

### v2

@@ 4.1 HardNet-Aff: Imposing Input-Dependent Affine Constraints
where $A(x)\in\mathbb{R}^{n_{\text{ineq}}\times n_{\text{out}}}$, $b(x)\in\mathbb{R}^{n_{\text{ineq}}}$, $C(x)\in\mathbb{R}^{n_{\text{eq}}\times n_{\text{out}}}$, $d(x)\in\mathbb{R}^{n_{\text{eq}}}$ for $n_{\text{ineq}}$ inequality and $n_{\text{eq}}$ equality constraints. For partitions $A(x)=[A_{(:n_{\text{eq}})}\;A_{(n_{\text{eq}}:)}]$ and $C(x)=[C_{(:n_{\text{eq}})}\;C_{(n_{\text{eq}}:)}]$, we make the following assumptions about the constraints:
@@
For each $x\in\mathcal{X}$, i) there exists $y\in\mathbb{R}^{n_{\text{out}}}$ that satisfies the constraints in (4),  ii) $C_{(:n_{\text{eq}})}$ is invertible, and iii) $\tilde{A}(x):=A_{(n_{\text{eq}}:)}\!-\!A_{(:n_{\text{eq}})}C_{(:n_{\text{eq}})}^{-1}C_{(n_{\text{eq}}:)}$ has full row rank.
Given $x\in\mathcal{X}$, when $C(x)$ has full row rank (i.e., no redundant constraints), there exists an invertible submatrix of $C(x)$ with its $n_{\text{eq}}$ columns. Without loss of generality, we can assume $C_{(:n_{\text{eq}})}$ is such submatrix by considering a proper permutation of the components of $f$. Then, the second assumption holds when the same permutation lets $C_{(:n_{\text{eq}})}$ invertible for all $x\in\mathcal{X}$. The last assumption requires the total number of the constraints $n_{\text{ineq}}+n_{\text{eq}}$ to be less than or equal to the output dimension $n_{\text{out}}$ and $A(x)$ to have full row rank. Also, note that handling the equality constraints separately, instead of viewing them as pairs of inequality constraints $C(x)f(x)\!\leq\!d(x)$ and $-C(x)f(x)\!\leq\!-d(x)$, is more efficient, because it decreases the total number of constraints from $n_{\text{ineq}}+2n_{\text{eq}}$ to $n_{\text{ineq}}+n_{\text{eq}}$ and avoids rank-deficient aggregated coefficient $[A(x);C(x);-C(x)]$.
Under the assumptions, we first efficiently reduce the $n_{\text{ineq}}+n_{\text{eq}}$ constraints to $n_{\text{ineq}}$ equivalent inequality constraints on partial outputs $f_{(n_{\text{eq}}:)}$ for a partition of the function $f(x)=[f_{(:n_{\text{eq}})};f_{(n_{\text{eq}}:)}]$. Consider the hyperplane in the codomain $\mathcal{Y}$ over which the equality constraints are satisfied. Then, for the function output $f(x)$ to be on the hyperplane, the first part $f_{(:n_{\text{eq}})}$ is determined by $f_{(n_{\text{eq}}:)}$:
@@
where $f_{\theta}^{*}(x):=f_{\theta}(x)-\tilde{A}(x)^{+}\mathsf{ReLU}{}\big{(}\tilde{A}(x)f_{\theta}(x)-\tilde{b}(x)\big{)}$ for all $x\in\mathcal{X}$ with $M^{+}:=M^{\top}(MM^{\top})^{-1}$ denoting the pseudoinverse of a matrix $M$. This novel projection satisfies the following properties; see Appendix A.1 for a proof.
@@
where $\bar{f_{\theta}}(x):=\big{[}[C_{(:n_{\text{eq}})}^{-1}\big{(}d(x)-C_{(n_{\text{eq}}:)}f_{\theta}(x)\big{)}];\;f_{\theta}(x)\big{]}\in\mathbb{R}^{n_{\textup{out}}}$.
Remark 4.3 (Parallel Projection).
Note that HardNet-Aff in (7) for $(n_{\text{ineq}},n_{\text{eq}})=(1,0)$ corresponds to the single-constraint case in (3). However, unlike (3), HardNet-Aff in general does not perform the minimum $\ell^{2}$-norm projection as in (2). Instead, its projection can be understood geometrically via Proposition 4.2 (iii). First, $f_{\theta}(x)$ is augmented to $\bar{f}_{\theta}(x)$ to satisfy the equality constraints. If $\bar{f}_{\theta}(x)$ violates any inequality constraints, it is projected onto the boundary of the feasible set ($a_{i}^{\top}\mathcal{P}(f_{\theta})(x)=b_{(i)}(x)$). Notably, the projection alters the output in a direction parallel to the boundary of each satisfied constraint’s feasible set ($a_{i}^{\top}\mathcal{P}(f_{\theta})(x)=a_{i}^{\top}\bar{f_{\theta}}(x)$), as shown in Fig. 2.
Figure 2: Illustration of input-dependent constraints and projections performed by HardNet. A target function $f\!:\!\mathbb{R}\!\rightarrow\!\mathbb{R}^{2}$ satisfies hard constraints $f(x)\!\in\!\mathcal{C}(x)$ for each $x\!\in\!\mathbb{R}$. The feasible set $\mathcal{C}(x)$ is visualized as the gray area for two sample inputs $x_{1}$ and $x_{2}$. While the function $f_{\theta}$ closely approximates $f$, it violates the constraints. HardNet-Aff projects the violated output onto the feasible set in parallel to the boundaries of the satisfied constraints. In contrast, the minimum $\ell^{2}$-norm optimization in (1) projects the output orthogonally to the closest boundary.
Remark 4.4 (Gradient Properties).
It is worth noting that, for $f_{\theta}(x)$ initialized outside the feasible set, the enforcement layer can yield zero gradients under certain conditions, even when the projected output differs from the target. However, such cases are infrequent in practice and can be mitigated as gradients are averaged over batched data. Thus, HardNet-Aff allows training the projected function to achieve values (strictly) within the feasible set using conventional gradient-based algorithms. Additionally, we can promote the model $f_{\theta}$ to be initialized within the feasible set using the warm-start scheme outlined in Appendix A.5, which involves training the model without the enforcement layer for a few initial epochs while regularizing constraint violations. Further details and discussions are provided in Appendix A.4.
While HardNet-Aff guarantees the satisfaction of the hard constraints (4), it should not lose the neural network’s expressivity for practical deployment. To that end, we rigorously show that HardNet-Aff preserves the neural network’s expressive power by the following universal approximation theorem; see Appendix A.2 for a proof.
@@ 4.2 HardNet-Cvx: Imposing General Input-Dependent Convex Constraints
Going beyond affine constraints, we discuss HardNet-Cvx as a conceptual framework for enforcing general input-dependent convex constraints $f(x)\in\mathcal{C}(x)\;\forall x\in\mathcal{X}$ where $\mathcal{C}(x)\subset\mathbb{R}^{n_{\text{out}}}$ is a convex set. Unlike the affine case, the closed-form projection from the single-constraint case in (3) does not extend to general convex constraints. Thus, we present HardNet-Cvx by generalizing the optimization-based projection in (2) as below:
@@
for all $x\in\mathcal{X}$. While no general closed-form solution for this optimization problem exists, we can employ differentiable convex optimization solvers for an implementation of HardNet-Cvx such as Amos & Kolter [5] for affine constraints (when HardNet-Aff cannot be applied) and Agrawal et al. [1] for more general convex constraints. This idea was briefly mentioned by Donti et al. [20] and used as a baseline (for input-independent constraints) by Tordesillas et al. [63]. However, the computational complexity of such solvers can be a limiting factor in time-sensitive applications. That said, we present HardNet-Cvx as a general framework, independent of specific implementation methods, to complement HardNet-Aff and provide a unified solution for various constraint types.
@@
Consider input-dependent constrained sets $\mathcal{C}(x)\subset\mathbb{R}^{n_{\textup{out}}}$ that are convex for all $x\in\mathcal{X}\subset\mathbb{R}^{n_{\textup{in}}}$. For any $p\in[1,\infty)$, let =$\mathcal{F}=\{f\in L^{p}(\mathcal{X},\mathbb{R}^{n_{\textup{out}}})|f(x)\in\mathcal{C}(x)\;\forall x\in\mathcal{X}\}$. Then, HardNet-Cvx with $w$-width $\mathsf{ReLU}{}$ neural networks defined in (8) universally approximates $\mathcal{F}$ if $w\geq\max\{n_{\textup{in}}+1,n_{\textup{out}}\}$.
@@ 5 Experiments
In this section, we demonstrate the versatility and effectiveness of HardNet-Aff over three constrained scenarios: learning with piecewise constraints, learning optimization solvers, and optimizing control policies in safety-critical systems. An additional scenario of learning safe decision logic for aircraft systems is introduced in Appendix A.8.
As evaluation metrics, we measure the violation of constraints in addition to the application-specific performance metrics. For a test sample $x\!\in\!\mathcal{X}$ and $n_{\text{ineq}}$ inequality constraints $g_{x}\big{(}f(x)\big{)}\!\leq\!0\!\in\!\mathbb{R}^{n_{\text{ineq}}}$, their violation is measured with the maximum ($\leq$ max) and mean ($\leq$ mean) of $\mathsf{ReLU}{}(g_{x}(f(x)))$ and the number of violated constraints ($\leq$ #). Similar quantities of $|h_{x}(f(x))|$ are measured for $n_{\text{eq}}$ equality constraints $h_{x}(f(x))=0\in\mathbb{R}^{n_{\text{eq}}}$. Then, they are averaged over all test samples. The inference time ($T_{\text{test}}$) for the test set and the training time ($T_{\text{train}}$) are also compared.
We compare HardNet with the following baselines: (i) NN: Plain neural networks, (ii) Soft: Soft-constrained neural networks, where constraint violations are penalized by adding regularization terms $\lambda_{\leq}\|\mathsf{ReLU}{}(g_{x}(f(x_{s})))\|_{2}^{2}+\lambda_{=}\|h_{x}(f(x_{s}))\|_{2}^{2}$ to the loss function for each sample $x_{s}\in\mathcal{X}$, (iii) DC3 [20]: Similarly to HardNet-Aff, DC3 approximates part of the target function with a neural network. It first augments the output to satisfy equality constraints, then corrects it using gradient descent to minimize inequality violations. DC3 backpropagates through this iterative correction procedure to train the model. All methods use 3-layer fully connected neural networks with 200 neurons per hidden layer and $\mathsf{ReLU}{}$ activation.
Figure 3: Learned functions at the initial (left) and final (right) epochs with the piecewise constraints. The models are trained on the samples indicated with circles, with their MSE from the true function shown in parentheses. HardNet-Aff adheres to the constraints from the start of the training and generalizes better than the baselines on unseen data. On the other hand, the baselines violate the constraints throughout the training.
Table 2: Results for the function fitting experiment. HardNet-Aff generalizes better than the baselines with the smallest MSE from the true function without any constraint violation. The max, mean, and the number of constraint violations are computed out of 401 test samples. Violations are highlighted in red. Standard deviations over 5 runs are shown in parentheses.
@@
NN 1.85 (0.19)
3.16/0.60/212.2 (0.29/0.04/7.19)
1.7 (0.01) 1.97 (0.02)
Soft 2.36 (0.22)
3.70/0.69/205.2 (0.25/0.03/1.94)
1.64 (0.08) 2.13 (0.12)
DC3 1.15 (0.10)
2.31/0.43/207.0 (0.23/0.02/3.58)
38.69 (0.33) 14.14 (0.47)
HardNet-Aff 0.16 (0.01) 0/0/0 (0/0/0) 6.38 (0.67) 4.48 (0.16)
5.1 Learning with Piecewise Constraints
In this experiment, we demonstrate the efficacy of HardNet-Aff and the expressive power of input-dependent constraints on a problem involving learning a function $f:[-2,2]\rightarrow\mathbb{R}$ with piecewise constraints shown in Fig. 3. The function outputs are required to avoid specific regions defined over separate subsets of the domain $[-2,2]$. An arbitrary number of such piecewise constraints can be captured by a single input-dependent affine constraint. The models are trained on 50 labeled data points randomly sampled from $[-1.2,1.2]$; see Appendix A.6 for details.
As shown in Fig. 3 and Table 2, HardNet-Aff consistently satisfies the hard constraints throughout training and achieves better generalization than the baselines, which violate these constraints. Especially at the boundaries $x\!=\!-1$ and $x=1$ in the initial epoch results, the jumps in DC3’s output value, caused by DC3’s correction process, insufficiently reduce the constraint violations. The performance of its iterative correction process heavily depends on the number of gradient descent steps and the step size. DC3 requires careful hyperparameter tuning unlike HardNet-Aff.
@@ 5.2 Learning Optimization Solver
Table 3: Results for learning solvers of nonconvex optimization problems with 100 variables, 50 equality constraints, and 50 inequality constraints. HardNet-Aff attain feasible solutions with the smallest suboptimality gap among the feasible methods. The max, mean, and the number of violations are computed out of the 50 constraints. Violations are highlighted in red. Standard deviations over 5 runs are shown in parentheses.
@@
Optimizer
-14.28 (0.00)
0/0/0 (0/0/0) 0/0/0 (0/0/0)
988.8 (5.74)
NN
-27.42 (0.00)
11.83/1.07/11.99 (0.03/0.00/0.01)
14.91/6.32/50 (0.01/0.00/0)
5.68 (1.52)
35.59 (1.58)
Soft
-12.08 (0.04)
0.00/0.00/0.00 (0.00/0.00/0.00)
0.46/0.17/49.98 (0.00/0.00/0.00)
6.34 (1.59)
35.51 (1.40)
DC3
-12.96 (0.05)
0.00/0.00/0.00 (0.00/0.00/0.00)
0.00/0.00/0.00 (0.00/0.00/0.00)
87.67 (2.40)
619.7 (9.16)
HardNet-Aff
-13.79 (0.13)
0/0/0 (0/0/0) 0/0/0 (0/0/0)
25.58 (4.76)
189.3 (1.95)
We consider the problem of learning optimization solvers as in Donti et al. [20] with the following nonconvex optimization problem:
$f(x)=\operatorname*{arg\,min}_{y}\;\;\frac{1}{2}y^{\top}Qy+p^{\top}\sin{y}\;\text{ s.t. }Ay\leq b,\;Cy=x\end{aligned},$ (9)
@@
In this experiment, we guarantee that the given constraints are feasible for all $x\in[-1,1]^{n_{\text{eq}}}$ by computing a proper $b$ for randomly generated $A,C$ as described in Donti et al. [20]. Then the models are trained on 10000 unlabeled data points uniformly sampled from $[-1,1]^{n_{\text{eq}}}$. To perform this unsupervised learning task, the loss function for each sample $x_{s}$ is set as $\frac{1}{2}f_{\theta}(x_{s})^{\top}Qf_{\theta}(x_{s})+p^{\top}\sin{f_{\theta}(x_{s})}$. To reproduce similar results as in Donti et al. [20], the models are equipped with additional batch normalization and dropout layers in this experiment. As shown in Table 3, HardNet-Aff consistently finds feasible solutions with a small suboptimality gap from the optimizer (IPOPT) with a much shorter inference time.
5.3 Optimizing Control Policy in Safety-Critical System
Table 4: Results for optimizing safe control policies. HardNet-Aff generates trajectories without constraint violation and has the smallest costs among the methods with zero violation. The max and mean constraint violations are computed for the violations accumulated throughout the trajectories. Violations are highlighted in red. Standard deviations over 5 runs are shown in parentheses.